OpenStudy (anonymous):

For the differential equation s" + bs' +8s=0, find the values of b that make the general solution overdamped, underdamped, or critically damped. I've gotten a few solutions, but the program is saying that my first interval for each is incorrect. My solutions: Overdamped: (-infinity,-4sqrt(2)),(4sqrt(2),infinity) Underdamped: (-4sqrt(2),0),(0,4sqrt(2)) Critically Damped: [-4sqrt(2),-4sqrt(2)],[4sqrt(2),4sqrt(2)]

OpenStudy (freckles):

ok so... \[a s'' +b s'+c s=0 \\ \text{ characteristic equation is } \\ ar^2+br+c=0 \\ \text{ so we have an overdamping if } b^2>4ac \\ \text{ we have an underdamping if } b^2<4ac \\ \text{ we have a critical damping if } b^2=4ac\] is this what you used?

OpenStudy (freckles):

\[\text{ overdamping if } b^2>32 \\ \text{ underdamping if } b^2<32 \\ \text{ critical damping if } b^2=32\]

OpenStudy (freckles):

\[\text{ Let } k \text{ be a constant greater than } 0 \\ b^2>a \text{ means } b \in (-\infty,-\sqrt{a}) \cup (\sqrt{a},\infty) \\ b^2<a \text{ means } b \in (-\sqrt{a},\sqrt{a}) \\ b^2=a \text{ means } b \in \{ -\sqrt{a}, \sqrt{a} \}\]

OpenStudy (anonymous):

I believe so yes, I may have messed up on the first intervals though

OpenStudy (freckles):

oops all of those a's were suppose to be k's

OpenStudy (freckles):

because I didn't want to use a for two different things :p

OpenStudy (anonymous):

I got you haha, let me see if maybe it would only be on interval on the last two

OpenStudy (anonymous):

One*

OpenStudy (freckles):

like the last one only has two solution so I didn't choose interval style to represent my solutions

OpenStudy (anonymous):

This is what I'm encountering

OpenStudy (anonymous):

The last two ended up being only one interval... maybe the first one is only one?

OpenStudy (anonymous):

Ah okay, they were all only one interval for some strange reason!

OpenStudy (anonymous):

Thanks for your help haha

OpenStudy (freckles):

one interval for each ?

OpenStudy (freckles):

the second one I thought should be one interval

OpenStudy (freckles):

(-4 sqrt(2),4 sqrt(2))

OpenStudy (freckles):

but the other two look fine to me given your instructions

OpenStudy (anonymous):

Yes, so the first one was (4sqrt(2),infinity), second was (0,4sqrt(2)), and third was [4sqrt(2),4sqrt(2)]

OpenStudy (anonymous):

For some reason, they just ended up being positive values.

OpenStudy (freckles):

yeah it looks like they ignored the negative values ...

OpenStudy (anonymous):

Yep, well I guess it's easier... weird that they didn't specify that though.

OpenStudy (freckles):

I'm going to do some more research...but that thing about that I posted about the overdamping b^2>4ac and so on... thing came from an MIT notes site

OpenStudy (anonymous):

Yeah, I remember seeing that in lecture and it definitely seems to work, it's just pulled from the quadratic equation haha... the part inside the square root!

OpenStudy (freckles):

right

OpenStudy (anonymous):

Thanks man, I appreciate your help

OpenStudy (anonymous):

haha nice